OpenStudy (jellybot23):

Identify the extrema for the function. Classify each as a local or absolute max or min value: f(x)=4X^3+15X^2-8X-3 How does one do this?

OpenStudy (freckles):

First step is to differentiate.

OpenStudy (jellybot23):

What do you mean?

OpenStudy (freckles):

Hmmm... so this isn't a calculus question?

OpenStudy (jellybot23):

It's pre-calc haha. I know what the graph should look like, but don't know how to algebraically get to the max and min values.

OpenStudy (freckles):

Well the only way I know how to do it without calculator is to differentiate. But you don't know how to differentiate so I don't know.

OpenStudy (jellybot23):

I just looked it up just to make sure I wasn't just having a brain freeze, but yeah we haven't learned that method :/

OpenStudy (jellybot23):

Thank you though

OpenStudy (freckles):

So the only other way I know is to graph it on your calculator and trace the graph to find the max(s) and min(s).

OpenStudy (freckles):

are you sure you never differentiated before? like finding f' given f(x)=2x^3 is f'(x)=2*3x^2 by constant multiple and power rule

OpenStudy (freckles):

You can also differentiate f by using the definition of derivative

OpenStudy (jellybot23):

Nope haha, not yet at least! Yeah when I plug it into my calculator, it looks like the min value is at -4 when X=1/3, but it's hard to tell, so I need an exact answer. I've done the derivative before

OpenStudy (freckles):

wait what?

OpenStudy (freckles):

you said nope and then the last sentence you say yes

OpenStudy (freckles):

that is contradicting information :p

OpenStudy (jellybot23):

Sorry I meant "I've not found..."

OpenStudy (jellybot23):

"I've not done the derivative before"

OpenStudy (jellybot23):

I'm looking up how to though

OpenStudy (freckles):

there is a long way and some short cut ways which the long way proves

OpenStudy (freckles):

teacher usually explains the long way first since it is the definition and also where the short cuts come from but then the class usually starts to use just the short cut ways but remembers the definition of derivative just in case it should come up somewhere

OpenStudy (freckles):

this is how you find the derivative of f: \[\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}=f'(x)\]

OpenStudy (freckles):

that is the definition of derivative

OpenStudy (freckles):

\[\small{f'(x)=\lim_{h \rightarrow 0} \frac{[4(x+h)^3+15(x+h)^2-8(x+h)-3]-[4x^3+15x^2-8x-3]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{4[(x+h)^3-x^3]+15[(x+h)^2-x^2]-8[(x+h)-x]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{4[(x+h-x)((x+h)^2+(x+h)(x)+x^2)+15(x+h-x)(x+h+x)-8h}{h}} \\ f'(x)=\lim_{h \rightarrow 0} \frac{4h[(x+h)^2+x(x+h)+x^2]+15h(2x+h)-8h}{h} \\ \text{ then recall } \frac{h}{h}=1 \\ \text{ so we have } \\ f'(x)= \lim_{h \rightarrow 0} \frac{4[(x+h)^2+x(x+h)+x^2]+15(2x+h)-8}{1} \frac{h}{h} \\ f'(x)= \lim_{ h \rightarrow 0} (4[(x+h)^2+x(x+h)+x^2]+15( 2x+h)-8 )\] now since the bottom isn't going to be 0 we can finally plug in 0 since this function is continuous at h=0 \[f'(x)=4[(x+0)^2+x(x+0)+x^2]+15(2x+0)-8 \\ f'(x)=4[3x^2]+15(2x)-8 \\ f'(x)=12x^2+30x-8\] but like I said there is a short cut way to do this then you find the critical numbers if there are any max(s) or min(s) they will occur at the critical points you can find the critical points for this one just by solving f'(x)=0 or that is solving 12x^2+30x-8=0

OpenStudy (jellybot23):

Oh my goodness, thank you for typing all of that out for me! I am going to write it down to refer to it when we do start learning it!

OpenStudy (freckles):

now the x values are where the max(s) and min(s) occur but f will give you the max(s) and min(s) once you find the critical numbers and plug those into f

OpenStudy (freckles):

if you are taking precal I'm almost certain you guys will cover the definition of derivative i'm surprised you haven't given this question

OpenStudy (freckles):

the way I went about it wasn't the only way to find the limit (or the derivative) I could have expanded things and cancel but instead I decided to group the cubes together and the squares together and linears together and factor the difference of cubes and factor the difference of squares

OpenStudy (freckles):

basically I used these two formulas: \[a^3-b^3=(a-b)(a^2+ab+b^2) \\ \text{ and } a^2-b^2=(a-b)(a+b)\]

OpenStudy (jellybot23):

Oh okay, haha it seems so intense! Thank you, for taking the time. I am writing it down, and I am going to ask my teacher about it :)

OpenStudy (freckles):

It will be intense and even may be confusing but with enough practice I believe you can get it and not think it so intimidating.

OpenStudy (jellybot23):

Yeah, I will definitely look over it and I'm going to check the lesson plan and see if we are learning it soon (I'm virtual schooled, so I can see what we are doing for the next few months)

OpenStudy (freckles):

by the way the local max should be closer to the point (-2.7,49.3) and the local min should be closer to the point (.2,-4) these aren't exact values but you can find the exact values by using that quadratic equation I gave you and inputting those values into f to find the corresponding y value

OpenStudy (freckles):

this is a cubic so there is no absolute extrema

OpenStudy (freckles):

only local

OpenStudy (jellybot23):

Okay, thank you :) I wonder if my teacher just expects us to guesstimate since we never learned that equation. And okay thank you! So you can tell from the degree if there are absolutes or locals?

OpenStudy (freckles):

since we were given no restrictions on the domain we normally assume all real numbers and yes you can look at the degree to help you determine the end behavior of the graph and this is what will help you know if there is a possibly absolute extrema

OpenStudy (jellybot23):

Alrighty! I know the behaviors of graphs, so that's a plus haha! Thank you so much for everything

OpenStudy (freckles):

np

OpenStudy (freckles):

I don't know how much I helped considering I didn't do it the way you guys were suppose to whatever way that was suppose to be... I kinda think since you guys haven't actually learned the derivative, it is just a guessimation thing by observing and tracing the graph with your ti83 (or whatever equivalent you have to that)

OpenStudy (jellybot23):

Yeah, I am going to guesstimate for now and then when I turn it in, I will ask her and mention the derivative :)